OpenStudy (anonymous):

Find the indefinite integral :

OpenStudy (anonymous):

\[\int\limits \ln (e ^{2x-1}) dx\]

OpenStudy (badhi):

$$\ln(e^{2x-1})=(2x-1)\ln(e)=(2x-1).1=(2x-1)$$